OpenStudy (anonymous):

Is the simplified form of 2 square root of 3 • 2 square root of 6 rational?

OpenStudy (whpalmer4):

Can you simplify it so all of the square root signs go away? If so, it is rational.

OpenStudy (anonymous):

I am trying to see how that is done. I think you do something like 2.2 radical 3.6, which turns into 4 radical 18, but now I don't know where to go? @whpalmer4

OpenStudy (anonymous):

seems unlikely \[2\sqrt3\times 2\sqrt6=4\sqrt{18}=4\sqrt{9\times 2}=12\sqrt2\] still got a radical there